OpenStudy (anonymous):

Evaluate the function at the specified value(s) of the independent variable and simplify. f(x)= -1/6x if x> -4 f(x)= 5-3x^2 if x≤ -4 Evaluate f(1) and f(-4)

OpenStudy (anonymous):

The above function is known as a "piecewise function." A piecewise function is a function that has two or more parts. To answer your question: to solve for f(1): 1 > -4, so f(1) = -1 / (6 * 1) = -1/6 to solve for f(-4): \[-4 \le -4\] so f(-4) = 5 - 3*(-4)^2 = 5 - 3 * 16 = 5 - 48 = -43 Hope I helped!